• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. C. Chang

    Log entry time 16:07:14 on June22,2005

    Entry number 146956

    keyword=Shift summary

    08:00 Shift started having problem with daq (ROC3 problem). Bob went dwon to all to cycle the rack power and that seems to fix the problem.
    08:45 Both left and right Q1 tripped due to low helium level. Call tech. Also call MCC to make a hall access since the tech needs to go down to the hall to reset power supplies.
    08:50 Also notice that right septum is also tripped.
    09:30 All magnets (left and right Q1 and right septum) are back.
    10:10 Call MCC to find out the status of the beam and was told that they are replacing one of the shunt to the magnet. It might take another 10 to 15 minutes before the beam is restored.
    11:00 MCC called to send us beam.
    15:25 Right septum tripped. Stop run.
    15:58 Call MCC to restore beam and resume data taking



    Run# Start Stop #Events Comments
    4819 08:46 11:01 229K Cosmic run
    4820 Junk
    4821 11:03 11:47 1M First good run of the shift
    4822 11:47 12:30 1M
    4823 12:30 13:12 1M
    4824 13:13 13:54 1M
    4825 13:54 14:12 370K Stopped due to ROC7 problem
    4826 14:16 15:00 1M
    4827 15:01 15:21 393K Right septum trip
    4828 15:57 On going